The rms current in an RL circuit is 0.21 A when it is connected to an ac generator with a frequency of 50 Hz and an rms voltage of 30 V .

Part A

Given that the inductor has an inductance of 170 mH , what is the resistance of the resistor?

Part B

Find the rms voltage across the resistor.

Part C

Find the rms voltage across the inductor.
Answer #1

(a) Z = Vrms / Irms
Z = 30 / 0.21
Z = 143 Ω

Z = √(R2 + ω2L2)
143 = √(R^2 + (2π * 50)^2*(0.170)^2)
R = 133 Ω

(b) Vrms,R = IrmsR
Vrms = (0.21) (132.65)
Vrms = 27.85 V = 28 V (rounded off)

(c) Vrms,L = IrmsωL
Vrms = (0.21)(2π * 50)(0.170)
Vrms = 11.21 V